Solve the system. If the system has infinitely many solutions, write the solution set with x arbitrary.
6x+5y=7
12x+10y=5

Answer by Alan3354(69443)

6x+5y=7
12x+10y=5
==========
12x+10y=14 Eqn 1 times 2
12x+10y=5
------------- Subtract
0 = 9
--> inconsistent, no solution